Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 78     Entry time: 01 Jun 2025, 08:48
Author: Yoonhee Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1400 uJ photon flux

Optical laser delivery

  • ns laser for particle illumination inside the chamber
  • another ns laser to visualize the tip of capillary on top of the chamber

Achievements/Observations

  • Measured GV for ~2 hours, hit rate is ~0.0003 (from Karabo TST/SFX/MONITOR/sfx, 1 = 100 %)
  • Measured PSI with different condition as below :
    1) without dryer : as we observed sample deposited on the tube connected to the inlet of dryer yesterday. Black tube is coated to avoid any deposition of particle on the wall)
    2) without both lasers : to avoid any heat effect if laser produces heat
    3) without EtOH : EtOH can disolve PSI crystals

Issues

  • PSI sample is not delivered well to the interaction region.
  • When adjust the skimmer at first, I did not run liquid flow. After the skimmer distance is adjusted to the reference pressure on Pressure at injector position, we started run sample and it increased pressures not only on pressure next to the skimmer and injector inlet, but also the pressure of IRU chamber. So it triggered the valve at upstream of IRU chamber (SPB_IRU_VAC/VALVE/VALVE_INLINE_1) to close. It is solved by adjusting skimmer gap again while the liquid flow is running.

Instrument status

  • AGIPD HV off.
  • AGIPD gate valve closed and pinned.
  • Attenuators set to thickest diamond.
  • Vacuum valves closed.
  • AGIPD breaks enguaged
  • ns Lasers off
  • HIREX is out
ELOG V3.1.4-7c3fd00